# Hypencoder: Hypernetworks for Information Retrieval

### v1

@@ 3.2. Query and Document Encoders
The Hypencoder framework is generic and can be applied to any implementation of query and document encoders. In this work, we use pretrained transformer-based encoder models commonly used in the recent neural network literature. Specifically, we use a pretrained BERT base model (Devlin et al., 2019) for encoding queries and documents. Even though Hypencoder can operate on all token representations produced for each document this work focuses on a single vector representation of documents, which is more efficient in terms of query latency, memory requirements, and disk usage. To do so, we can either use the contextualized embedding representing the [CLS] token or take the mean of all the contextualized embeddings for all the non-pad input tokens. Empirically, we found that using the [CLS] token performs better. Therefore, the document representation produced by the encoder is a single vector with 768 dimensions, i.e., the same as BERT’s output dimensionality. We refer to it as $E_{d}\in\mathbb{R}^{m\times h}$, where $m=1$ in our setting.
@@ 3.3. The Hyperhead Layers
For improved clarity, we focus only on the weight creation process as the biases are created in the exact same way. The contextualized query embeddings $E_{q}\in\mathbb{R}^{n\times h}$ produced by the query encoder are independently transformed by $l$ hyperhead layers, each of which corresponds to a layer in the q-net. Each hyperhead layer converts the embeddings $E_{q}$ into key and value matrices:
(2) $K^{q}_{i}=\theta_{K_{i}}\times[E_{q};1]\quad\quad V^{q}_{i}=\theta_{V_{i}}\times[E_{q};1]$
where $\theta_{K_{i}},\theta_{V_{i}}\in\mathbb{R}^{h\times h}$ denote learnable parameters for constructing key and value matrices. In the above equation, the embedding matrix $E_{q}$ is concatenated with a column of all ones (i.e., $[E_{q};1]$) to model both weight multiplication and bias addition.
Each key matrix $K_{i}$ and value matrix $V_{i}$ will be used for the creation of the weights in the $i$th layer of the q-net. With the keys and values in hand, single-head scaled-dot-product attention (Vaswani et al., 2017) is performed using a query matrix $Q_{i}\in\mathbb{R}^{r\times h}$ where $r$ is the layer dimensionality in the $i$th layer of q-net. In our case, all of the weights except the last layer are square matrices, making $r=h$. Each $Q_{i}$ is a set of learnable embeddings, similar to those used as input tokens for transformer models. Hence, the hidden layer representation $H_{i}\in\mathbb{R}^{r\times h}$ is then computed as follows:
@@
A ReLU activation (Agarap, 2018) is then applied to each $H_{i}$ followed by layer normalization (Ba et al., 2016). Next a point-wise feed-forward layer is applied to produce $\widehat{H}^{q}_{i}$:
@@
where L-Norm denotes layer normalization. Note that each weight in q-net has a unique $\theta_{W_{i}}$ and $\theta_{b_{i}}$. There are no learnable parameters in layer normalization.
@@
where $\theta_{H_{i}}\in\mathbb{R}^{r\times h}$ is the same size as $\widehat{H}^{q}_{i}$ and acts as a base weight which allows the model to learn universal (i.e., query-independent) patterns that are applicable for all queries.
The process for the bias vectors is identical except the query matrix used in the attention operation $Q_{i}\in\mathbb{R}^{r\times h}$ has $r=1$ as there is only a single column in the output.
@@ 3.4. The q-net Network
Weights and biases produced by the hyperhead layers are not by themselves a neural network. They need a certain arrangement and additional components (e.g. non-linearity). This is where the Hypencoder’s q-net converter comes in. The converter knows the architecture of the q-net and given the weights and biases from the hyperhead layers, it produces a callable neural network object which takes as input the document representation $E_{d}$.
@@
In our experiments, we use a simple feed-forward architecture for the q-net. The output $x^{d}_{i+1}$ for the input $x^{d}_{i}$ at a given layer $i$ is given by:
(6) $x^{d}_{i+1}=\text{L-Norm}\left(\text{ReLU}(W^{q}_{i}(x^{d}_{i})+b^{q}_{i})\right)+x^{d}_{i}$
where L-Norm represents a layer normalization without learnable parameters and the addition of $x^{d}_{i}$ is a residual connection. No residual connection is applied before the final layer (i.e., layer $l$). The layer in Equation (6) is repeated $l$ times. Finally, a relevance score is produced using a linear projection layer with an output dimensionality of 1.
@@ 3.5. Training
Training Hypencoder is no different from training a bi-encoder as it shares the same core components, i.e., a query encoder and document encoder. The only difference is instead of using an inner product to find the similarity the q-net is applied to the document representations. Thus, our contributions are solely the architecture and not a specific training technique. In this paper we employ a simple distillation training setup, for more details see Section 4.2
@@ 3.6. Efficient Retrieval using Hypencoder
To study the extent to which Hypencoder’s retrieval can be approximated for efficient retrieval, we developed an approximate search technique based loosely on navigating small world graphs (Lifshits and Zhang, 2009; Malkov and Yashunin, 2020). In the index stage we construct a graph where documents are nodes connected to their neighbors by edges. We use $L_{2}$ distance between document embeddings similar to (Tan et al., 2020).
@@
Algorithm 1 Hypencoder Efficient Search
@@ 3.7. Comparison to Existing Neural IR Methods
We argue that Hypencoder can exactly reproduce existing neural ranking models. Let us start by formalizing the main components of existing neural methods: (1) a query representation $E_{q}$, (2) a document representation $E_{d}$, (3) some combination function $f_{c}(\cdot,\cdot)$, and (4) the final neural network that produces a score $f_{s}(\cdot)$. In comparison, Hypencoder does not have an $f_{c}(\cdot,\cdot)$ as the q-net takes $E_{d}$ as its only input.
We will now demonstrate that Hypencoder can exactly replicate any neural retrieval method that has the components above. The first step is to include $E_{q}$ and $f_{c}(\cdot,\cdot)$ in the q-net. This allows the q-net to exactly produce the input to $f_{s}(\cdot)$ that the existing neural retriever used. Next we reproduce the neural function from $f_{s}(\cdot)$ with query-dependent weights in the q-net. When $E_{d}$ is input to the q-net, all the original components of the existing neural retrieval model are present and thus the score can be exactly replicated. There is one difference, which is the weights of $f_{s}(\cdot)$ are query-dependent. However, this can be remedied in two ways: (1) shared weights can be used for all queries exactly replicating the original neural method (2) the weights for $f_{c}$ can have a common non-query-dependent base weight, similar to our implementation (see details in Section 3.3), this way if there is no benefit using query-dependent weights the shared weight can be used, but if there is additional benefit the model and optimizer can learn to take advantage of it. Thus, Hypencoder can not only exactly replicate all existing neural retrieval methods it also allows the model to dynamically leverage query-dependent weights when the model determines they are beneficial.
@@ 4.1.1. Training Dataset
The dataset used for training our models is the training split of the MSMARCO passage retrieval dataset (Nguyen et al., 2016) which contains 8.8M passages and has 503K training queries with at least one corresponding relevant passage. The queries in the MSMARCO training set are short natural language questions asked by users of the Bing search engine.
@@ 4.1.3. Evaluation Datasets
For in-domain performance, we use the MSMARCO Dev set (Nguyen et al., 2016), TREC Deep Learning 2019 (Craswell et al., 2020b), and TREC Deep Learning 2020 (Craswell et al., 2020a). The MSMARCO Dev set contains around 7k queries with shallow labels, the majority of queries only have a single passage labeled as relevant. This collection uses queries from the same distribution as the training queries making it a clear test of the in-domain performance. On this dataset we report the standard evaluation metrics: MRR and Recall@1000. The TREC Deep Learning 2019 and 2020 datasets have a similar query distribution to MSMARCO Dev but feature far fewer queries, i.e., 97 queries combined. The lower number of queries is compensated by far deeper annotations with every query having several annotated passages.
@@
Figure 2. Relationship between the three main parameters of our efficient search: the size of the initial set of candidates $\tilde{C}$, the number of neighbors to explore $nCanddidates$, and the number of iterations $maxIter$ and both effectiveness in terms of TREC DL ’19 nDCG@10 and efficiency in terms of Query Latency.
To assess out-of-domain performance, we evaluate on question answering tasks on different domains, specifically, the TREC COVID (Roberts et al., 2021) and NFCorpus datasets (Boteva et al., 2016) for the biomedical domain and FiQA (Maia et al., 2018) for the financial domain. We also evaluate on DBPedia (Hasibi et al., 2017) as an entity retrieval dataset and on Touché (Bondarenko et al., 2020) as an argument retrieval dataset. We use the BEIR (Thakur et al., 2021) versions of these datasets from the ir_datasets library. [^3] Available at https://ir-datasets.com/.
To explore the full capabilities of Hypencoder we want to evaluate how it performs on retrieval tasks that are more challenging than standard question-passage retrieval tasks. To some extent hardness is subjective, but we tried to define a clear set of criteria to define difficulty: (1) current neural retrieval models should struggle on the task, (2) term matching models like BM25 should also struggle on the task, (3) the queries are longer or otherwise more complicated than standard web queries. An additional requirement we had was that for tasks that were significantly different from the MSMARCO training data we wanted adequate training data to fine-tune the models before evaluation. We believe this is reasonable as we are not investigating the models’ zero-shot performance but the inherent limits of the model.
The first dataset we select was the TREC Tip-of-the-Tongue (TOT) 2023 (Arguello et al., 2023) that contains queries written by users that know many aspects of the item they are looking for but not the name of the item. Thus TOT queries tend to be verbose and can include many facets. The TREC TOT 2023 dataset specifically looks at TOT queries for movies with the corresponding movie’s Wikipedia page as the golden passage. We use the development set as the test set relevance labels are not public yet. There are 150 queries. Each query has a single relevant passage. For training we use the data from Bhargav et al. (2022) which is around 15k TOT queries from Reddit for the book and movie domain.
@@ 4.2.1. Training Details
All the Hypencoders use BERT (Devlin et al., 2019) base uncased as the base model. We use PyTorch and Huggingface for model implementation and training. All of our q-nets use a input dimension of 768 and hidden dimension of 768. Unless otherwise stated, we use 6 linear layers in the model not including the final output projection layer.
We use a training batch size of 64 per device and 128 in total. A single example in the batch is a query, positive document, and 8 additional documents ranging in relevance. The positive document is the top ranked document by our teacher model. The other 8 documents are sampled randomly from the passages associated with the query. For more details about the dataset see Section 4.1.1. Passages were truncated to 196 tokens and queries to 32 tokens.
Our primary loss function is Margin MSE (Hofstätter et al., 2020). When computing the loss, we construct (query, positive document, negative document) triplets where all of the negatives for a query form their own triplet. The loss is found by averaging the individual loss of all the triplets. In addition to Margin MSE, we use an in-batch cross entropy loss where the (query, positive document) is assumed to be a true positive and all the other queries’ positive documents are assumed to be negatives. We do not consider the additional “hard” negatives from the query in the cross entropy loss as many of these documents are relevant to the query. We use AdamW as our optimizer with default settings and a learning rate of 2e-5 with a constant scheduler after a warm up of 6k steps. Our training hardware consist of two A100 GPUs with 80GB memory. Training took around 6 days.
To select the best model we evaluate each model on the validation set every 50k steps and pick the model with the best R Precision within the first 800k steps. We selected R Precision due to the fact that it balances both recall and precision in a single metric and does not require a predefined cutoff.
When training for the harder tasks we use AdamW with the learning rate 8e-6 with a linear scheduler and a warm-up ratio of 1e-1. For TOT training we train for 25 epochs or 3.3k steps. For FollowIR training we train for 1 epoch or around 10k steps. We use a batch-size of 96 and cross entropy loss. Each example in the batch includes a query, positive document, and hard negative document. We use a maximum document and query length of 512 tokens.
@@ 4.3. Baselines
For comparison with Hypencoder, we include several baseline models and models which we include for reference which are not directly comparable. Our main baselines which we evaluate on all datasets are TAS-B (Hofstätter et al., 2021), CL-DRD (Zeng et al., 2022), BM25, and our own bi-encoder baseline which we call BE-Base. We train BE-Base exactly the same as Hypencoder except we use separate encoders and use a linear LR scheduler. We select our main dense baselines TAS-B and CL-DRD as they are both strong bi-encoder models which leverage knowledge-distillation training and which use the same document embedding dimension of 768. For in-domain results we include an additional set of dense models: ANCE (Xiong et al., 2021), TCT-ColBERT (Lin et al., 2020), and MarginMSE (Hofstätter et al., 2020). We only include these models in the in-domain results to save space in other sections and because TAS-B and CL-DRD outperform the other baselines. For reference we also include: the late-interaction model ColBERT v2 (Santhanam et al., 2022); the neural sparse model SPLADE++ SD (Formal et al., 2022); RepLLaMA a 7b parameter bi-encoder model (Ma et al., 2024); DRAGON a bi-encoder trained with 5 teacher models and 20M synthetic queries; MonoBERT a reranking model reranking the top 1k BM25 retrieavls (Nogueira et al., 2019); and the reranking model cross-SimLM reranking the top 200 passages from bi-SimLM (Wang et al., 2023). Reference results were taken from the RepLLaMA (Ma et al., 2024) and DRAGON (Lin et al., 2023a) papers.

### v2

@@ 3.2. Query and Document Encoders
The Hypencoder framework is generic and can be applied to any implementation of query and document encoders. In this work, we use pretrained transformer-based encoder models commonly used in the recent neural network literature. Specifically, we use a pretrained BERT base model (Devlin et al., 2019) for encoding queries and documents. Although Hypencoder can operate on all document token representations this work focuses on a single vector representation of documents, which is more efficient in terms of query latency, memory requirements, and disk usage. To do so, we can either use the contextualized embedding representing the [CLS] token or take the mean of all the contextualized embeddings for all the non-pad input tokens. Empirically, we find that using the [CLS] token performs better. Therefore, the document representation produced by the encoder is a single vector with 768 dimensions, i.e., the same as BERT’s output dimensionality. We refer to it as $E_{d}\in\mathbb{R}^{m\times h}$, where $m=1$ in our setting.
@@ 3.3. The Hyperhead Layers
In the finalized implementation, the contextualized query embeddings $E_{q}\in\mathbb{R}^{n\times h}$, produced by the query encoder, are independently transformed by $l$ hyperhead layers where $l$ is the number of weight matrices and bias vectors needed by the $q$-net. For improved clarity, we focus only on the weight creation process for $W_{i}^{q}\in\mathbb{R}^{r\times t}$ which is the weight matrix for the $i$th layer of the $q$-net. The values of $r$ and $t$ are selected based on the desired shape of the final weight matrix. The first step is to turn $E_{q}$ into key and value matrices:
(2) $K^{q}_{i}=[E_{q};1]\times\theta_{K_{i}}\quad\quad V^{q}_{i}=[E_{q};1]\times\theta_{V_{i}}$
where $\theta_{K_{i}},\theta_{V_{i}}\in\mathbb{R}^{(h+1)\times t}$ denote learnable parameters for constructing key and value matrices. In the above equation, the embedding matrix $E_{q}$ is concatenated with a column of all ones (i.e., $[E_{q};1]$) to model both weight multiplication and bias addition.
With the keys and values in hand, single-head scaled-dot-product attention (Vaswani et al., 2017) is performed using a query matrix $Q_{i}\in\mathbb{R}^{r\times t}$. The matrix $Q_{i}$ is a set of learnable embeddings, similar to those used as input tokens for transformer models. Hence, the hidden layer representation $H_{i}\in\mathbb{R}^{r\times t}$ is computed as follows:
@@
A ReLU activation (Agarap, 2018) is then applied to each $H_{i}$ followed by layer normalization (Ba et al., 2016) along the last dimension. Next a point-wise feed-forward layer is applied to produce $\widehat{H}^{q}_{i}\in\mathbb{R}^{r\times t}$:
@@
where L-Norm denotes layer normalization and $\theta_{W_{i}}\in\mathbb{R}^{t\times t}$ and $\theta_{b_{i}}\in\mathbb{R}^{t\times 1}$ are learnable parameters. Note, there are no learnable parameters in the layer normalization.
@@
where $\theta_{H_{i}}\in\mathbb{R}^{r\times t}$ is the same size as $\widehat{H}^{q}_{i}$ and acts as a base weight which allows the model to learn universal (i.e., query-independent) patterns that are applicable for all queries.
The process for the bias vectors is identical except the query matrix used in the attention operation $Q_{i}\in\mathbb{R}^{r\times t}$ has $r=1$ as there is only a single column in the output.
@@ 3.4. The $q$-net Network
The weights and biases produced by the hyperhead layers are not by themselves a neural network. They need a certain arrangement and additional components (e.g. non-linearity). This is where the Hypencoder’s $q$-net converter comes in. The converter knows the architecture of the $q$-net and given the weights and biases from the hyperhead layers, it produces a callable neural network object which takes as input the document representation $E_{d}$.
@@
We use a simple feed-forward architecture for the $q$-net. The output $x^{d}_{i+1}$ for the input $x^{d}_{i}$ at a given layer $i$ is given by:
(6) $x^{d}_{i+1}=\text{L-Norm}\left(\text{ReLU}(W^{q}_{i}(x^{d}_{i})+b^{q}_{i})\right)+x^{d}_{i},$
where L-Norm represents a layer normalization without learnable parameters and the addition of $x^{d}_{i}$ is a residual connection. No residual connection is applied before the final layer. The layer in Equation (6) is repeated until the last layer where a relevance score is produced using a linear projection layer with an output dimensionality of 1.
@@ 3.5. Training
Training Hypencoder is no different from training a bi-encoder as it shares the same core components, i.e., a query encoder and document encoder. The only difference is that instead of using an inner product to find the similarity, the $q$-net is applied to the document representations. Thus, our novel contributions are focused solely on the architecture and not on a specific training technique. In this paper we employ a simple distillation training setup, for more details see Section 4.2.
@@ 3.6. Efficient Retrieval using Hypencoder
To study the extent to which Hypencoder’s retrieval can be approximated for efficient retrieval, we developed an approximate search technique based loosely on navigating small world graphs (Lifshits and Zhang, 2009; Malkov and Yashunin, 2020). In the indexing stage we construct a graph where documents are nodes connected to their neighbors by edges. We use $\ell^{2}$ distance between document embeddings similar to Tan et al. (Tan et al., 2020).
@@
Algorithm 1 Hypencoder Efficient Search.
@@ 3.7. Comparison to Existing Neural IR Methods
We argue that Hypencoder can exactly reproduce existing neural ranking models. Let us start by formalizing the main components of existing neural methods: (1) a query representation $E_{q}$, (2) a document representation $E_{d}$, (3) some combination function $f_{c}(\cdot,\cdot)$, and (4) the final neural network that produces a score $f_{s}(\cdot)$. In comparison, Hypencoder does not have an $f_{c}(\cdot,\cdot)$ as the $q$-net takes $E_{d}$ as its only input. We will now demonstrate that Hypencoder can exactly replicate any neural retrieval method that has the components above. The first step is to include $E_{q}$ and $f_{c}(\cdot,\cdot)$ in the $q$-net. This allows the $q$-net to exactly produce the input to $f_{s}(\cdot)$ that the existing neural retriever used. Next we reproduce the neural function from $f_{s}(\cdot)$ with query-dependent weights in the $q$-net. When $E_{d}$ is input to the $q$-net, all the original components of the existing neural retrieval model are present and thus the score can be exactly replicated. There is one difference, which is the weights of $f_{s}(\cdot)$ are query-dependent. However, this can be remedied in two ways: (1) shared weights can be used for all queries exactly replicating the original neural method (2) the weights for $f_{c}$ can have a common query-independent base weight, similar to our implementation (see details in Section 3.3), this way if query-dependent weights are not beneficial the shared weight can be used, but if there are additional benefits the model and optimizer can learn to take advantage of them. Thus, Hypencoder can not only exactly replicate all existing neural retrieval methods it also allows the model to dynamically leverage query-dependent weights when the model determines they are beneficial.
@@ 4.1.1. Training Dataset
For training our models we used the MSMARCO passage retrieval dataset (Nguyen et al., 2016) which contains 8.8M passages and has 503K training queries with at least one relevant passage. The queries in the MSMARCO training set are short natural language questions asked by users of the Bing search engine.
@@ 4.1.3. Evaluation Datasets
For in-domain performance, we use the MSMARCO Dev set (Nguyen et al., 2016), TREC Deep Learning 2019 (Craswell et al., 2020b), and TREC Deep Learning 2020 (Craswell et al., 2020a). The MSMARCO Dev set contains around 7k queries with shallow labels, the majority of queries only have a single passage labeled as relevant. This collection uses queries from the same distribution as the training queries making it a clear test of the in-domain performance. On this dataset we report the standard evaluation metrics: MRR and Recall@1000. The TREC Deep Learning 2019 and 2020 datasets have a similar query distribution to MSMARCO Dev but feature far fewer queries (97 queries combined). The lower number of queries is compensated by far deeper annotations with every query having several annotated passages.
@@
Figure 2. Relationship between the three main parameters of our efficient search: the size of the initial set of candidates $\tilde{C}$, the number of neighbors to explore $nCanddidates$, and the number of iterations $maxIter$ and both effectiveness in terms of TREC DL ’19 nDCG@10 and efficiency in terms of query latency in milliseconds.
To assess out-of-domain performance, we evaluated question-answering tasks from various domains. Specifically, the TREC COVID (Roberts et al., 2021) and NFCorpus datasets (Boteva et al., 2016) for the biomedical domain and FiQA (Maia et al., 2018) for the financial domain. We also evaluated on DBPedia (Hasibi et al., 2017) as an entity retrieval dataset and on Touché (Bondarenko et al., 2020) as an argument retrieval dataset. We use the BEIR (Thakur et al., 2021) versions of these datasets from the ir_datasets library. [^3] Available at https://ir-datasets.com/.
To explore the full capabilities of Hypencoder we want to evaluate how it performs on retrieval tasks that are more challenging than standard question-passage retrieval tasks. To some extent hardness is subjective, but we tried to define a clear set of criteria to define difficulty: (1) current neural retrieval models should struggle on the task, (2) term matching models like BM25 should also struggle on the task, (3) the queries are longer or otherwise more complicated than standard web queries. An additional requirement we had for tasks that were significantly different from the MSMARCO was training data to facilitate fine-tuning prior to evaluation. We believe this is reasonable as we are not investigating the models’ zero-shot performance but the inherent limits of the model.
The first hard dataset is the TREC Tip-of-the-Tongue (TOT) Track from 2023 (Arguello et al., 2023). It contains queries written by users that know many aspects of the item they are looking for but not the name of the item. Thus, TOT queries tend to be verbose and can include many facets. The TREC TOT 2023 dataset specifically looks at TOT queries for movies with the corresponding movie’s Wikipedia page as the golden passage. We used the 150 query development set as the test set relevance labels were not public at the time of writing. Each query has a single relevant passage. For training we use the data from Bhargav et al. (2022) which is around 15k TOT queries from Reddit for the book and movie domains.
@@ 4.2.1. Training Details
All the Hypencoders use BERT (Devlin et al., 2019) base uncased as the base model. We use PyTorch and Huggingface Transformers for model implementation and training. All of our $q$-nets use an input dimension of 768 and a hidden dimension of 768. Unless otherwise stated, we use 6 linear layers in the $q$-net not including the final output projection layer.
We used a training batch size of 64 per device and 128 in total. A single example in the batch is a query, positive document, and 8 additional documents ranging in relevance. The positive document is the top ranked document by our teacher model. The other 8 documents are sampled randomly from the passages associated with the query. For more details about the dataset see Section 4.1.1. Passages were truncated to 196 tokens and queries to 32 tokens.
Our primary loss function is Margin MSE (Hofstätter et al., 2020). We construct (query, positive document, negative document) triplets where all of the negatives for a query form their own triplet. The loss is found by averaging the individual loss of all the triplets. In addition to Margin MSE, we use an in-batch cross entropy loss where the (query, positive document) is assumed to be a true positive and all the other queries’ positive documents are assumed to be negatives. We do not consider the additional “hard” negatives from the query in the cross entropy loss as many of these documents are relevant to the query. We use AdamW as our optimizer with default settings and a learning rate of 2e-5 with a constant scheduler after a warm up of 6k steps. Our training hardware consists of two A100 GPUs with 80GB memory. Training took around 6 days.
To select the best model, we evaluate each model on the validation set every 50k steps and pick the model with the best R Precision within the first 800k steps. We selected R Precision because it balances both recall and precision in a single metric and does not require a predefined cutoff.
When training for the harder tasks we use AdamW with the learning rate 8e-6 with a linear scheduler and a warm-up ratio of 1e-1. We forgo Margin MSE and only use cross entropy loss. For TOT training we train for 25 epochs or 3.3k steps. For FollowIR training we train for 1 epoch or around 10k steps. We use a batch-size of 96. Each example in the batch includes a query, positive document, and hard negative document. We use a maximum document and query length of 512 tokens.
@@ 4.3. Baselines
For comparison with Hypencoder, we include several baseline models and models which we include for reference which are not directly comparable (see 4.4.1 for more details). Our main baselines which we evaluate on all datasets are TAS-B (Hofstätter et al., 2021), CL-DRD (Zeng et al., 2022), BM25, and our own bi-encoder baseline which we call BE-Base. We train BE-Base exactly the same as Hypencoder except we use separate encoders and use a linear LR scheduler. We select our main dense baselines TAS-B and CL-DRD as they are both strong bi-encoder models which leverage knowledge-distillation training and which use the same document embedding dimension of 768. For in-domain results we include an additional set of dense models: ANCE (Xiong et al., 2021), TCT-ColBERT (Lin et al., 2020), and MarginMSE (Hofstätter et al., 2020). We only include these models in the in-domain results to save space in other sections and because TAS-B and CL-DRD outperform the other baselines. For reference we also include: the late-interaction model ColBERT v2 (Santhanam et al., 2022); the neural sparse model SPLADE++ SD (Formal et al., 2022); RepLLaMA a 7b parameter bi-encoder model (Ma et al., 2024); DRAGON a bi-encoder trained with 5 teacher models and 20M synthetic queries; MonoBERT a reranking model reranking the top 1k BM25 retrieavls (Nogueira et al., 2019); and the reranking model cross-SimLM reranking the top 200 passages from bi-SimLM (Wang et al., 2023). Reference results were taken from the RepLLaMA (Ma et al., 2024) and DRAGON (Lin et al., 2023a) papers.